If log{(a+b)/2}=1/2(loga+logb) Prove that a=b Please give the solution

  • If log{(a+b)/2}=1/2(loga+logb)
  • Prove that a=b
  • Please give the solution

Grade:11

1 Answers

Rinkoo Gupta
9 years ago
log{(a+b)/2}=1/2 log(ab)
=log(ab)1/2
(a+b)/2= (ab)1/2
a+b=2(ab)1/2
(va)2 +(vb )2-2vavb=0

(va -vb)2=0

va-vb=0

va=vb

a=b

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty